OpenStudy (anonymous):

Solve the DE: (3x-y)dx + (x+y)dy = 0 I've tried it using integrating factors, but (dM/dy - dN/dx)/N = -2/(x+y) & (dN/dy - dM/dx)/M = 0

OpenStudy (anonymous):

ops i meant to say (dM/dy - dN/dx)/N = -2/(x+y) & (dN/dx - dM/dy)/M = 0

OpenStudy (anonymous):

just noticed that (dN/dx - dM/dy)/M = 2/(3x-y)

OpenStudy (anonymous):

im trying this problem with a different aproach, here is what i got:\[-(3x-y)/(x+y)= dy/dx\]\[0= (dy/dx)(x+y)+3-y\] Im kind of lost regarding of what to do next.

OpenStudy (unklerhaukus):

can you arrange the origional DE in to the form dy/dx = f (y/x)

OpenStudy (unklerhaukus):

\[\frac{dy}{dx}=\frac{-(3x-y)}{x+y}\\\qquad=-\frac{3x/x-y/x}{x/x+y/x}\\=\]

OpenStudy (anonymous):

I still don't see how i could solve\[\frac{ 3-y/x }{ 1+y/x }\]

OpenStudy (unklerhaukus):

try the subsitution y/x = v y = vx dy = ...+...

OpenStudy (anonymous):

oooo, i think i can work from there on, thanks for the help.

OpenStudy (unklerhaukus):

** dy/dx = ... + ...

OpenStudy (unklerhaukus):

have you found dy/dx?

OpenStudy (dumbcow):

this is not an exact diff equ ... you can't explicitly solve for "y" in terms of "x"

OpenStudy (dumbcow):

look at solution http://www.wolframalpha.com/input/?i=%283x-y%29dx+%2B+%28x%2By%29dy+%3D+0